OpenStudy (anonymous):

If A standard electrical outlet provides a 120-V,60Hz sinewave. Would the average voltage be 108.08616?if the peak to peak voltage is 339.36

OpenStudy (anonymous):

if the peak to peak voltage is 339.36

OpenStudy (anonymous):

Do you mean: a standard outlet provides 120V AC with 60Hz ?

OpenStudy (anonymous):

yeah i forget that

OpenStudy (theeric):

The average voltage does not depend on the frequency! http://www.electronics-tutorials.ws/blog/average-voltage.html \(\dfrac{1}{2}V_\text{pk-pk}=V_\text{pk}\) So we can use the equation at that link to find that \(V_\text {ave}=\dfrac{2}{\pi}V_\text{pk}=\dfrac{1}{\pi}V_\text{pk-pk}\)